OpenStudy (anonymous):

Indicate whether the following statement is true or false: A linear function eventually exceeds a quadratic function with a positive leading coefficient.

OpenStudy (anonymous):

@HWBUSTER00 can you help?

OpenStudy (anonymous):

Consider y = 5x and y = 5x². Which one would increase faster?

OpenStudy (anonymous):

y=5^2 @peachpi

OpenStudy (anonymous):

yes. So the quadratic exceeds the linear

OpenStudy (anonymous):

so its true @peachpi

OpenStudy (anonymous):

no it's false. your question is the other way around.

OpenStudy (anonymous):

ohh. mind helping me with another question? @peachpi

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The first-, second-, and third-year enrollment values for a technical school are shown in the table below. Enrollment at a Technical School Year (x) First Year f(x) Second Year s(x) Third Year t(x) 2009 785 756 756 2010 740 785 740 2011 690 710 781 2012 732 732 710 2013 781 755 800 Which of the following statements is true based on the data in the table? The solution to f(x) = s(x) is x = 2,009. The solution to f(x) = s(x) is x = 785. The solution to f(x) = t(x) is x = 2,010. The solution to f(x) = t(x) is x = 740.